# Valid F-screening in linear regression

### v1

@@ 3 Computing selective p-values in a retrospective analysis
${p_{H_{0}^{M}\mid E}}$ $=\Pr\Bigg{(}W/Z\geq F_{H_{0}^{M}}\cdot\frac{m}{n-p-1}$
$\Big{|}W-cZ\geq\mathrm{RSE}^{2}\cdot\left(F_{H_{0}^{M}}\cdot m-\frac{\mathrm{R}^{2}}{1-\mathrm{R}^{2}}\cdot(n-p-1)\right)/\sigma^{2}\Bigg{)}.$ (19)
Furthermore, approximating ${\tilde{\sigma}^{2}}$ in (16) via Monte Carlo only requires knowledge of $n$, $p$, $\alpha_{0}$, and $Y^{\top}(I_{n-1}-P_{X})Y=\mathrm{RSE}^{2}\cdot(n-p-1)$; therefore, ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}$ (17) can also be computed retrospectively.
In contrast, confidence interval and point estimate computations that account for Step 1 of F-screening (described in Appendix B) rely on statistics that are not generally reported in software output and cannot be easily obtained retrospectively.
@@ 4 The special case of ANOVA
F-screening is particularly pervasive in the context of ANOVA, where it dates back almost a century (Fisher, 1935; Lindquist, 1940). As a simple example, suppose that we are interested in how plant height, $Y$, varies between treatments by fertilizers $A$, $B$, and $C$. We fit the model
$Y=\beta_{A}\cdot\bm{I}_{A}+\beta_{B}\cdot\bm{I}_{B}+\beta_{C}\cdot\bm{I}_{C}+\epsilon,$ (20)
where $\bm{I}_{A}$, $\bm{I}_{B}$, and $\bm{I}_{C}$ are $n$-dimensional indicator variables for treatment with fertilizers $A$, $B$, and $C$. In the language of ANOVA, $H_{0}^{1:p}:\beta_{A}=\beta_{B}=\beta_{C}$ is often referred to as the “omnibus test.” If $H_{0}^{1:p}$ is rejected, then one typically conducts post hoc tests (Everitt & Skrondal, 2010) to see how the expected plant height differs across fertilizers; for instance, we may test $H_{0}:\beta_{A}=\beta_{B}$. This post hoc test can be equivalently expressed as a test of $H_{0}:\beta_{B\hskip 1.42262pt\text{vs.}\hskip 1.42262ptA}=0$ in the reparameterized model
@@
The machinery developed in Sections 23 for valid inference in Step 2 of F-screening applies directly in the setting of ANOVA. In fact, Section 3 simplifies a bit, since in the context of ANOVA, it is not necessary to have access to $F_{H_{0}^{M}}$, $\mathrm{R}^{2}$, and $\mathrm{RSE}$ to conduct a retrospective analysis: instead, it suffices to know the number of samples within each group, and the mean and standard deviation of the outcome variable within each group. This is made precise in the following remark.
@@
Consider a one-way ANOVA with $K$ groups, parameterized as in (4). Suppose we have access to only the following quantities, for $k=1,\ldots,K$:
@@
[…]
where we define $n=\sum_{k=1}^{K}n_{k}$, $\bar{y}=\frac{1}{n}\sum_{k=1}^{K}\hat{\mu}_{k}n_{k}$, $\text{SSB}=\sum_{k=1}^{K}n_{k}(\hat{\mu}_{k}-\bar{y})^{2}$, and $\text{SSW}=\sum_{k=1}^{K}(n_{k}-1)\hat{\sigma}_{k}^{2}$. Thus $p_{H_{0}^{1:K}}=1-F_{K-1,n-K}(F_{H_{0}^{1:K}})$, and ${p_{H_{0}^{M}\mid E}}$ can be computed via (19).
@@ 5.1 Type 1 error
We generate data according to (1) with $n=100$, $p=10$, $\sigma^{2}=1$, $X_{ij}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}\mathcal{N}(0,1)$, and $\beta_{1}=\beta_{2}=\ldots=\beta_{10}=0$, so that $H_{0}^{1:p}$ in (2) holds. We consider testing $H_{0}^{1}:\beta_{1}=0$ if and only if $H_{0}^{1:p}$ is rejected at level $\alpha_{0}=0.05$ using ${F_{H_{0}^{1:p}}}$ from (3) as the test statistic. For the subset of $100000$ simulated datasets for which $H_{0}^{1:p}$ is rejected, the left-hand panel of Fig. 4 displays (i) $p_{H_{0}^{M}}$ from (5), which does not account for rejection of $H_{0}^{1:p}$; (ii) ${p_{H_{0}^{M}\mid E}}$ from (12), which accounts for rejection of $H_{0}^{1:p}$ and requires known variance; (iii) ${p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}}$ from (15), which uses the standard variance estimator and therefore does not control the selective Type 1 error; and (iv) ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}$ from (17), which uses a conditionally debiased variance estimator. We see that ${p_{H_{0}^{M}\mid E}}$ controls selective Type 1 error, as does ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}$ when $\alpha$ is not large (i.e., in the regime of interest). Because $p_{H_{0}^{M}}$ and ${p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}}$ do not control the selective Type 1 error, we do not consider them in Section 5.2.
@@ 5.2 Power
First, we formalize the sample splitting procedure. We randomly partition $n$ independent realizations of (1) into two equally-sized sets: we use the first for training and the second for testing. We then use the training set to test $H_{0}^{1:p}$ using an F-statistic (which we refer to as $F_{H_{0}^{1:p}}^{\mathrm{train}}$), and, if and only if we reject $H_{0}^{1:p}$, we use the test set to test $H_{0}^{M}$ using an F-statistic, which yields the p-value $p_{H_{0}^{M}}^{\text{test}}$. Since the training and test sets are independent, $p_{H_{0}^{M}}^{\text{test}}$ is uniformly distributed under $H_{0}^{M}$, regardless of the value of $F_{H_{0}^{1:p}}^{\mathrm{train}}$.
Now we return to the simulation set-up. We generate data according to (1) with $n=100$, $p=10$, $\sigma^{2}=1$, $X_{ij}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}\mathcal{N}(0,1)$, $\beta_{2}=\ldots=\beta_{10}=0$, and with $\beta_{1}$ ranging from 00 to $0.8$.
First, we test $H_{0}^{1:p}$ (2) at $\alpha_{0}\in\left\{0.05,0.1,0.5\right\}$ using either (i) all the data, or (ii) just the training data. The middle panel of Fig. 4 displays the proportion of simulated datasets for which $H_{0}^{1:p}$ is rejected. Unsurprisingly, we see that we have substantially more power to reject $H_{0}^{1:p}$ using all the data versus using only the training data.
@@
Moreover, it is worth noting that the benefits of our selective p-value over the sample splitting p-value extend beyond the empirical results seen in Fig. 4. Depending on how sample splitting is performed, the testable hypotheses may differ between train and test sets (e.g., if rank deficiency emerges from random partitioning). Also, it may be hard, in an applied setting, to convince a collaborator to use only part of their data to test $H_{0}^{1:p}$ and to reserve the remainder for testing $H_{0}^{M}$, as is required for sample splitting. By contrast, our selective p-values enable the collaborator to use all of the data to test $H_{0}^{1:p}$, thereby yielding a result for the test of $H_{0}^{1:p}$ that coincides with the standard (uncorrected) F-screening procedure. Furthermore, the sample splitting p-values cannot be computed retrospectively using summary statistics, as in Sections 3 and 4.
@@ 5.3 Confidence intervals
In this subsection, we investigate confidence interval coverage and width. The computation of confidence intervals is described in Appendix B.
@@
To examine confidence interval width, we first vary $\beta_{1}\in\left[-1,1\right]$ while holding $n=100$ constant. We set $p=5$, $\sigma^{2}=1$, and $\beta_{2}=\dots=\beta_{5}=0$. For each fixed value of $\beta_{1}\in(0,1)$, we generate data according to (1) until we obtain 1000 datasets where the overall hypothesis test is rejected. For each dataset, we compute (i) the selective confidence interval using the true value of $\sigma^{2}$, and (ii) the selective confidence interval using the debiased $\tilde{\sigma}^{2}$. (We do not depict the standard confidence interval, as we showed in the left-hand panel of Fig. 5 that it is invalid.) Finally, we compute the average confidence interval width across the 1000 datasets and plot these values as a function of $\beta_{1}$ in the middle panel of Fig. 5. In the right-hand panel, we repeat the analysis in the middle panel, but with $\beta_{1}=0$ and $n\in\{50,\ldots,500\}.$
We now comment on the shape of the curves shown in the middle panel of Fig. 5. The selective confidence intervals—whether computed using the true value of $\sigma^{2}$ or the debiased estimate of $\sigma^{2}$—exhibit a striking shape: they are narrow both when $\beta_{1}\approx 0$ and when $|\beta_{1}|$ is large. This pattern can be understood by separately considering the cases where (i) $\left\lvert\beta_{1}\right\rvert$ is large, (ii) $\left\lvert\beta_{1}\right\rvert$ is moderate, and (iii) $\left\lvert\beta_{1}\right\rvert$ is small. When $\left\lvert\beta_{1}\right\rvert$ is large, the event $E_{1}$ (6) is highly likely, in which case conditional inference roughly coincides with standard inference, and so the intervals are no wider than the standard (non-selective) intervals. When $\left\lvert\beta_{1}\right\rvert$ is moderate, the event $E_{1}$ (6) is only moderately likely, and when it occurs, it often happens that the data falls near the boundary of the selection region, and little signal regarding $\beta_{1}$ remains for inference, leading to wide confidence intervals. Finally, when $\left\lvert\beta_{1}\right\rvert$ is small, the event $E_{1}$ (6) is unlikely to occur; however, when it does occur, selection is often driven by (chance) association between $X_{2},\ldots,X_{p}$ and $Y$, in which case sufficient signal about $\beta_{1}$ remains for inference, leading to moderately narrow confidence intervals.
@@ 6 Data analysis
In this section, we demonstrate a valid data analysis involving F-screening, both with and without access to the full data. First we re-analyze the data of Xu & Wu (2018), and report selective p-values, confidence intervals, and point estimates for each coefficient. Then we conduct a retrospective analysis of Keil et al. (2023), using only the standard ANOVA outputs, because we do not have access to their data: in this setting, only the selective p-values are reported. In both analyses, we compare the selective quantities to their standard counterparts.
@@ 6.1 Prospective inference
Xu & Wu (2018) examined the publicly-available National Health and Nutrition Examination Survey (NHANES) dataset from Centers for Disease Control and Prevention (2024) (CDC) in order to investigate possible changes in femoral neck bone mineral density (BMD)  [^4] BMD is coded as “DXXNKBMD” in the NHANES dataset. across the years 2005 to 2014 in US adults over the age of 30. In the NHANES dataset, years are binned into 2005–2006, 2007–2008, 2009–2010, 2013–2014; no data from 2011–2012 is available for BMD. We restrict our analysis to the $n=7135$ males in the dataset, and consider the model
@@
where $\bm{I}_{\text{5--6}}$, $\bm{I}_{\text{7--8}}$, $\bm{I}_{\text{9--10}}$, and $\bm{I}_{\text{13--14}}$ are indicator variables for the time periods 2005–2006, 2007–2008, 2009–2010, and 2013–2014, respectively. Note that (22) contains age, as reported in the NHANES dataset, as a covariate. For convenience, we regress age out of the response and the predictors, leading to the model
$U^{\top}\text{BMD}=\beta_{\text{5--6}}\cdot U^{\top}\bm{I}_{\text{5--6}}+\beta_{\text{7--8}}\cdot U^{\top}\bm{I}_{\text{7--8}}+\beta_{\text{9--10}}\cdot U^{\top}\bm{I}_{\text{9--10}}+\beta_{\text{13--14}}\cdot U^{\top}\bm{I}_{\text{13--14}}+U^{\top}\epsilon,$ (23)
where $U\in\mathbb{R}^{7135\times 7134}$ has orthonormal columns and satisfies $UU^{\top}:=I_{n}-P_{\text{AGE}}$. In this setting, the overall null hypothesis is
@@
and test $H_{0}^{\text{ 7--8 vs. 5--6}}:\beta_{\text{7--8 vs. 5--6}}=0.$ We take a similar approach for the other five pairwise tests. Results are shown in Table 1. In this example, accounting for the rejection of $H_{0}^{1:p}$ in (24) has virtually no impact—the selective and standard quantities are identical up to the third decimal.
@@
2005–06 versus 2007–08 -0.004 (-0.013, 0.005) 0.423 -0.004 (-0.013, 0.005) 0.423
2005–06 versus 2009–10 -0.003 (-0.012, 0.006) 0.473 -0.003 (-0.012, 0.006) 0.473
2005–06 versus 2013–14 -0.020 (-0.029, -0.010) $0$0.020 (-0.029, -0.010) 0.000
2007–08 versus 2009–10 0.000 (-0.008, 0.009) 0.923 0.000 (-0.008, 0.009) 0.923
2007–08 versus 2013–14 -0.016 (-0.025, -0.007) $0$0.016 (-0.025, -0.007) 0.000
2009–10 versus 2013–14 -0.017 (-0.025, -0.008) $0$0.017 -(0.025, -0.008) 0.000

### v2

@@ 3 Selective p-values in a retrospective analysis
${p_{H_{0}^{M}\mid E}}=\Pr\Bigg(\tfrac{W}{Z}\geq F_{H_{0}^{M}}\cdot\tfrac{m}{n-p-1}\Big|W-cZ\geq\tfrac{\mathrm{RSE}^{2}}{\sigma^{2}}\cdot\left(F_{H_{0}^{M}}\cdot m-\tfrac{\mathrm{R}^{2}}{1-\mathrm{R}^{2}}\cdot(n-p-1)\right)\Bigg).$ (23)
Approximating $\tilde{\sigma}^{2}$} (20) via Monte Carlo only requires knowledge of $n$, $p$, $\alpha_{0}$, and $y^{\top}(I_{n-1}-P_{X})y=\mathrm{RSE}^{2}\cdot(n-p-1)$; therefore, $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21) can also be computed retrospectively.
In contrast, the selective confidence interval and point estimates described in Appendix B involve quantities that are not generally reported in software output, and thus cannot be easily obtained retrospectively.
@@ 4 The special case of ANOVA
F-screening is particularly pervasive in the context of ANOVA, where it dates back almost a century (Fisher, 1935; Lindquist, 1940). As a simple example, suppose that we are interested in how plant height, $Y$, varies between treatments by fertilizers $A$, $B$, and $C$. We fit the model $Y=\beta_{A}\cdot\bm{I}_{A}+\beta_{B}\cdot\bm{I}_{B}+\beta_{C}\cdot\bm{I}_{C}+\epsilon,$ where $\bm{I}_{A}$, $\bm{I}_{B}$, and $\bm{I}_{C}$ are $n$-dimensional indicator variables for treatment with fertilizers $A$, $B$, and $C$. In the language of ANOVA, $H_{0}^{1:p}:\beta_{A}=\beta_{B}=\beta_{C}$ is often referred to as the “omnibus test.” If $H_{0}^{1:p}$ is rejected, then one typically conducts post hoc tests (Everitt & Skrondal, 2010) to see how the expected plant height differs across fertilizers; for instance, we may test $H_{0}:\beta_{A}=\beta_{B}$. This post hoc test can be equivalently expressed as a test of $H_{0}:\beta_{B\hskip 1.42262pt\text{vs.}\hskip 1.42262ptA}=0$ in the reparameterized model
@@
The machinery developed in Sections 23 for valid inference in Step 2 of F-screening applies directly in the setting of ANOVA. In fact, retrospective analysis, as detailed in Section 3, is even more straightforward because in the context of ANOVA, it is not necessary to have access to $F_{H_{0}^{M}}$, $\mathrm{R}^{2}$, and $\mathrm{RSE}$ to conduct a retrospective analysis: instead, it suffices to know the number of samples within each group, and the mean and standard deviation of the outcome variable within each group. This is made precise in the following remark.
@@
Consider a one-way ANOVA with $K$ groups, parameterized analogously to (4). Suppose we have access to only the following quantities, for $k=1,\ldots,K$:
@@
$\mathrm{R}^{2}=\tfrac{\text{SSB}}{\text{SSB}+\text{SSW}},\quad\mathrm{RSE}=\sqrt{\tfrac{\text{SSW}}{n-K}},\quad F_{H_{0}^{k_{1}\hskip 0.7113pt\text{vs.}\hskip 0.7113ptk_{2}}}=\tfrac{(n-K)(\hat{\mu}_{k_{1}}-\hat{\mu}_{k_{2}})^{2}}{\text{SSW}\left(\frac{1}{n_{k_{1}}}+\frac{1}{n_{k_{2}}}\right)},\quad F_{H_{0}^{1:K}}=\tfrac{\text{SSB}/(K-1)}{\text{SSW}/(n-K)},$
where we define $n=\sum_{k=1}^{K}n_{k}$, $\bar{y}=\frac{1}{n}\sum_{k=1}^{K}\hat{\mu}_{k}n_{k}$, $\text{SSB}=\sum_{k=1}^{K}n_{k}(\hat{\mu}_{k}-\bar{y})^{2}$, and $\text{SSW}=\sum_{k=1}^{K}(n_{k}-1)\hat{\sigma}_{k}^{2}$. Thus $p_{H_{0}^{1:K}}=1-F_{K-1,n-K}(F_{H_{0}^{1:K}})$, and $p_{H_{0}^{M}\mid E}$} can be computed via (23), and further we can approximate $\tilde{\sigma}^{2}$} and compute $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21).
@@ 5.1 Type 1 error
We generate data according to (1) with $n=100$, $p=10$, $\sigma^{2}=1$, $X_{ij}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}\mathcal{N}(0,1)$, and $\beta_{1}=\ldots=\beta_{10}=0$, so that $H_{0}^{1:p}$ in (2) holds. We consider testing $H_{0}^{1}:\beta_{1}=0$ if and only if $H_{0}^{1:p}$ is rejected at level $\alpha_{0}=0.05$ using $F_{H_{0}^{1:p}}$} (3). For the subset of $100000$ datasets for which $H_{0}^{1:p}$ is rejected, the left-hand panel of Fig. 5 displays (i) $p_{H_{0}^{M}}$ from (5), which does not account for rejection of $H_{0}^{1:p}$; (ii) $p_{H_{0}^{M}\mid E}$} from (10), which accounts for rejection of $H_{0}^{1:p}$ and requires known variance; (iii) $p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}$} from (19), which uses the standard variance estimator and does not control the selective Type 1 error; and (iv) $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} from (21), which uses a conditionally debiased variance estimator. We see that $p_{H_{0}^{M}\mid E}$} controls selective Type 1 error, as does $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} when $\alpha$ is not large (i.e., in the regime of interest). Because $p_{H_{0}^{M}}$ and $p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}$} do not control the selective Type 1 error, we do not consider them in Section 5.2.
@@ 5.2 Power
First, we formalize the sample splitting procedure. We randomly partition $n$ independent realizations of (1) into two equally-sized training and test sets. We use the training set to test $H_{0}^{1:p}$ using an F-statistic (which we refer to as $F_{H_{0}^{1:p}}^{\mathrm{train}}$), and, if and only if we reject $H_{0}^{1:p}$, we use the test set to test $H_{0}^{M}$ using an F-statistic, which yields the p-value $p_{H_{0}^{M}}^{\text{test}}$. Since the training and test sets are independent, $p_{H_{0}^{M}}^{\text{test}}$ is uniformly distributed under $H_{0}^{M}$, regardless of the value of $F_{H_{0}^{1:p}}^{\mathrm{train}}$.
Now we return to the simulation set-up. We generate data according to (1) with $n=100$, $p=10$, $\sigma^{2}=1$, $X_{ij}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}\mathcal{N}(0,1)$, $\beta_{2}=\ldots=\beta_{10}=0$, and with $\beta_{1}$ ranging from 0 to $0.8$.
First, we test $H_{0}^{1:p}$ (2) at $\alpha_{0}\in\left\{0.05,0.1,0.5\right\}$ using either (i) all the data, or (ii) just the training data. The middle panel of Fig. 5 displays the proportion of simulated datasets for which $H_{0}^{1:p}$ is rejected. Unsurprisingly, we have substantially more power to reject $H_{0}^{1:p}$ using all the data versus using only the training data.
@@
It is worth noting that the benefits of our selective p-value over the sample splitting p-value extend beyond the empirical results seen in Fig. 5. Depending on how sample splitting is performed, the testable hypotheses may differ between train and test sets (e.g., if rank deficiency emerges from random partitioning). Also, it may be hard in an applied setting to convince a collaborator to use only part of their data to test $H_{0}^{1:p}$ and to reserve the remainder for testing $H_{0}^{M}$. By contrast, our selective p-values enable the collaborator to use all of the data to test $H_{0}^{1:p}$, thereby yielding a result for the test of $H_{0}^{1:p}$ that coincides with the standard (uncorrected) F-screening procedure. Furthermore, the sample splitting p-values cannot be computed retrospectively using summary statistics, as in Section 3.
@@ 5.3 Confidence intervals
We now investigate coverage and width of the confidence intervals derived in Appendix B.
@@
To examine confidence interval width, we first vary $\beta_{1}\in\left[-1,1\right]$ while holding $n=100$ constant. We set $p=5$, $\sigma^{2}=1$, and $\beta_{2}=\dots=\beta_{5}=0$. For each fixed value of $\beta_{1}\in\left[-1,1\right]$, we generate data according to (1) until we obtain 1000 datasets where the overall hypothesis test is rejected. For each dataset, we compute (i) the selective confidence interval using the true value of $\sigma^{2}$, and (ii) the selective confidence interval using the debiased $\tilde{\sigma}^{2}$. (We do not depict the standard confidence interval, as we showed in the left-hand panel of Fig. 6 that it is invalid.) Finally, we compute the average confidence interval width across the 1000 datasets and plot these values as a function of $\beta_{1}$ in the middle panel of Fig. 6. In the right-hand panel, we repeat the analysis in the middle panel, but with $\beta_{1}=0$ and $n\in\{50,\ldots,500\}.$
We now comment on the shape of the curves shown in the middle panel of Fig. 6. The selective confidence intervals—whether computed using the true value of $\sigma^{2}$ or the debiased estimate of $\sigma^{2}$—exhibit a striking shape: they are narrow both when $\beta_{1}\approx 0$ and when $|\beta_{1}|$ is large. This pattern can be understood by separately considering the cases where (i) $\left\lvert\beta_{1}\right\rvert$ is large, (ii) $\left\lvert\beta_{1}\right\rvert$ is moderate, and (iii) $\left\lvert\beta_{1}\right\rvert$ is small. When $\left\lvert\beta_{1}\right\rvert$ is large, the event $E_{1}$ (6) is highly likely, in which case conditional inference roughly coincides with standard inference, and so the intervals are not appreciably wider than the standard (non-selective) intervals. When $\left\lvert\beta_{1}\right\rvert$ is moderate, the event $E_{1}$ (6) is only moderately likely, and when it occurs, it often happens that the data falls near the boundary of the selection region, and little signal regarding $\beta_{1}$ remains for inference, leading to wide confidence intervals. Finally, when $\left\lvert\beta_{1}\right\rvert$ is small, the event $E_{1}$ (6) is unlikely to occur; however, when it does occur, selection is often driven by (chance) association between $X_{2},\ldots,X_{p}$ and $Y$, in which case sufficient signal about $\beta_{1}$ remains for inference, leading to moderately narrow confidence intervals.
@@ 6 Data analysis
In this section, we demonstrate a valid data analysis involving F-screening, both with and without access to the full data. First we re-analyze the data of Xu & Wu (2018), and report selective p-values, confidence intervals, and point estimates for each coefficient. Then we conduct a retrospective analysis of Keil et al. (2023) and compute the selective p-values using only the standard ANOVA outputs because we do not have access to their data. In both analyses, we compare the selective quantities to their standard counterparts.
@@ 6.1 Prospective inference
Xu & Wu (2018) used the publicly-available National Health and Nutrition Examination Survey (NHANES) dataset from Centers for Disease Control and Prevention (2024) (CDC) to investigate possible changes in femoral neck bone mineral density (BMD)  [^4] BMD is coded as “DXXNKBMD” in the NHANES dataset. across the years 2005 to 2014 in US adults over the age of 30. In the NHANES dataset, years are binned into 2005–2006, 2007–2008, 2009–2010, 2013–2014; no data from 2011–2012 is available for BMD. We restrict our analysis to the $n=7135$ males in the dataset, and consider the model
@@
where $\bm{I}_{\text{5--6}}$, $\bm{I}_{\text{7--8}}$, $\bm{I}_{\text{9--10}}$, and $\bm{I}_{\text{13--14}}$ are indicator variables for the time periods 2005–2006, 2007–2008, 2009–2010, and 2013–2014, respectively. Note that (25) contains age, as reported in the NHANES dataset, as a covariate. For convenience, we regress out age: let $U\in\mathbb{R}^{7135\times 7134}$ such that $UU^{\top}:=I_{7135}-P_{\text{AGE}}$ and $U^{\top}U=I_{7135}$, and consider the model
$U^{\top}\text{BMD}=\beta_{\text{5--6}}\cdot U^{\top}\bm{I}_{\text{5--6}}+\beta_{\text{7--8}}\cdot U^{\top}\bm{I}_{\text{7--8}}+\beta_{\text{9--10}}\cdot U^{\top}\bm{I}_{\text{9--10}}+\beta_{\text{13--14}}\cdot U^{\top}\bm{I}_{\text{13--14}}+U^{\top}\epsilon.$ (26)
In this setting, the overall null hypothesis is
@@
and test $H_{0}^{\text{ 7--8 vs. 5--6}}:\beta_{\text{7--8 vs. 5--6}}=0.$ The other five pairwise tests are conducted similarly. Results are shown in Table 1. In this case, accounting for the rejection of $H_{0}^{1:p}$ (27) has virtually no impact—the selective and standard quantities are identical up three decimals.
@@
05–06 vs. 07–08 -0.004 (-0.013, 0.005) 0.423 -0.004 (-0.013, 0.005) 0.423
05–06 vs. 09–10 -0.003 (-0.012, 0.006) 0.473 -0.003 (-0.012, 0.006) 0.473
05–06 vs. 13–14 -0.020 (-0.029, -0.010) 0.000 -0.020 (-0.029, -0.010) 0.000
07–08 vs. 09–10 0.000 (-0.008, 0.009) 0.923 0.000 (-0.008, 0.009) 0.923
07–08 vs. 13–14 -0.016 (-0.025, -0.007) 0.000 -0.016 (-0.025, -0.007) 0.000
09–10 vs. 13–14 -0.017 (-0.025, -0.008) 0.000 -0.017 -(0.025, -0.008) 0.000